n-factor of reaction , FeS2+O2 -> Fe2O3+SO2..?


n-factor of FeS2?

Grade:11

5 Answers

Sunil Kumar FP
9 years ago
For acids, n-factor is defined as the number of H+ions replaced by 1 mole of acid in a reaction.
For bases, n-factor is defined as the number of OH–ions replaced by 1 mole of base in a reaction.
The n-factor for salts which react such that no atom undergoes redox reaction is defined as the total moles of catioinic/anionic charge replaced in 1 mole of the salt.
n factor for FeS2 is 1
because in ferrous sulphide the oxidation number changes from +2 to +3
Ritesh Pratap Singh
6 years ago
sorry, the answer u`ve given is wrong because here in FeS2,both Fe and S are oxidised. Fe from +2 to +3 and S from -1 to +4 therefore n factor of FeS2 is 11.
J Vijaya bhaskar Gowd
6 years ago
the balanced reaction will be 4FeS2+11O2   ->   2Fe2O3+8SO2
 
It is a redox reaction.Number of electrons gained =Number of electrons lost.
1 mole of FeS2 = (11/4) moles of O2
Each O atom takes 2 electrons for reduction. Therefore for 1 O2 molecule 4 electrons are absorbed for reduction.
Therefore number of moles of electrons gained by O2 = (11/4)*4=11.
Therefore number of moles of electrons lost by 1 mole of FeS2=11 moles of electrons.
THerefore equivalent Weight of FeS2 in this reaction = molecular weight/11
Kashish
6 years ago
            But how can you balance this reaction when the medium (acidic or basic) is not given? Please explain how balancing is done.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

For acids, n-factor is defined as the number of H+ions replaced by 1 mole of acid in a reaction.
For bases, n-factor is defined as the number of OH–ions replaced by 1 mole of base in a reaction.
The n-factor for salts which react such that no atom undergoes redox reaction is defined as the total moles of catioinic/anionic charge replaced in 1 mole of the salt.
Hence n factor for FeS2 is 1 because in ferrous sulphide the oxidation number changes from +2 to +3
Thanks and Regards